HUBBLE SPACE TELESCOPE DAILY REPORT #990 PERIOD COVERED: 0800 EST Daily Status Report as of 284/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 5059 (Seven Points of Light _ Cycle 2 Calibration - Fixed) The Fine Guidance Sensors were used to make observations of standard star UPGREEN-69. The objective was to calibrate the single star interferometeric fringe pattern at seven different points in the central region of the astrometric FGS. The astrometric observations were monitored and no problems were seen in the telemetry. 1.2 Completed Three sets of GHRS/2 4344 (The Mass of White Dwarf V471 Tuari - Cycle 3 High) The High Resolution Spectrograph was used to make observations of V471 Tau (Visual Magnitude 10.7), which is an eclipsing binary star consisting of a white dwarf and a K2 class main sequence star. It is hoped that the strong Lyman-Alpha absorption of the white dwarf can be used to determine its velocity curve, and that that can be used to calculate its mass. A total of eight observations were received and displayed. The onboard target acquisition was successful. Eight observations were received and displayed. The observations were reported to look normal. 1.3 Completed FGS 4892 (Duplicity Among Hyades Stars - Cycle 3 High) The FGSs were used to make astrometric observations of the possible binary star VA537 which is located in the Hyades open cluster. The OSS monitored the observation, and no anomalies were reported. 1.4 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make two high latitude medium-deep sky survey observations taken in parallel with the prime science observations. Both observations were received and displayed. They were reported to look okay. One recentering event occurred during the second observation. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed PC 5072 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary camera were used to make Earth Flat calibration exposures using the wide and medium band filters. Four observations were received and displayed in the OSS. Two observations were reported to be saturated, and the other two were normal. 1.7 Completed PC 3614 (Boron an a Probe of Stellar Structures and Mass Loss) The High Resolution Spectrograph was used to make four spectral observations of the Boron lines in star HD262462. The abundance of Boron can be used to study the process on the interior of these stars. All four spectral observations were reported to look okay. 1.8 Completed FGS 5100 (HIPPARCOS Plate Scale Calibration - Cycle 3 Instrument Scientist Calibration) Fine Guidance Sensor #3 was used to make POS MODE observations of stars HP39-2250 AND HP39-225-R. This pair of stars were observed also with the HIPPARCOS Astrometry Satellite. These observations will be used to do an overall average calibration of the plate scale on FGS #3. The astrometric observations were monitored in the OSS, and no problems were seen in the telemetry. No anomalies were reported. 1.9 Completed FOS/RD/BL 3791 (Quasar Absorption Line Survey: Cycle 2 Observations) FGS #3 was used to make astrometric observations of star HP-42-29-RE which was also observed with the HIPPARCOS astrometry satellite. These observations will be used to make an overall average calibration of the plate scale of FGS #3. There was at least one recentering event which occurred during the observations. 1.10 Completed Two Sets of PC 5023 (Calibration/WFPC Cycle 3 Polarization Calibration) The Wide Field Camera was used to make polarization calibration observations by using images of the non-polarized standard star PKS-52243-12. This will allow for the absolute transmittance of the polarized optical path to be established. Three images were received and displayed. All three images were reported to be okay. 1.11 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. The images were reported to look fine. 1.12 Completed Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. Ten observations were received and displayed. Four were reported to be saturated, and six were normal. 1.13 Completed HSP/POL 4887 (X- Ray Binaries - Augmentation - Cycle 3 High) The High Speed Photometer was used to make photometric and polarimetric observations of X-Ray binary star 4U1907+09 to obtain more information about the structure and condition in the gas streams connecting the primary star to its neutron companion. The observations were executed as scheduled. The onboard target acquisition was successful. 12 observations were received and displayed. All of their photometry observations were reported to look normal. 1.14 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of four observations were received. All four images were reported to be underexposed. 1.15 Completed Two Sets of PC 5073 (WF/PC Cycle 3 Calibration: PC Flats Wide. 2) The Planetary Camera was used to make two sets of Earth flat calibration images using different filter combinations. These images will be used for flat field corrections of science images made with the PC. Both images were reported to be underexposed. 1.16 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. Two observations were received and displayed. Both were reported to look normal. 1.17 Completed PC 5071 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 1) The Planetary Camera was used to make Earth calibration flat images which will be used to remove the pixel to pixel variations in gain of the chips. The flats from this proposal will be used for flat field corrections for subsequent science data. The observations were executed as scheduled. Eight observations were received and displayed. Four were underexposed, and four were okay. 1.18 Completed PC 4308 (Snapshot Survey For Companions to Planetary - Nebula Nuclei - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula IC289. It is possible that the central star in a double star. The observations was received and displayed in the OSS, the data was reported to be normal. 1.19 Completed HSP/POL 4942 (Active Galactic Nuclei - Cycle 2 High) The High Speed photometer was used to make observations of the active galactic nucleus of variable galactic object BL-Lac. This is to monitor the intensity and linear polarization of the radiation emitted by the active galactic nuclei. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were noted. 1.20 Completed FGS 5249 (HIPPARCOS Plate Scale Calibration: Part 2 Cycle 3 Instrument Scientist Calibration) FGS #3 was used to make POS Mode astrometric observations of standard star HP48-30345 which was also observed using the HIPPARCOS astrometry satellite. These observations will be used to do an overall average calibration of the plate scale of FGS #3. Most of the observation occurred during LOS. The section which was able to be monitored was reported to look normal, except for a single failure to acquire fine lock. 1.21 Completed FOC/96 3804 (Imaging of UV Bright Stars in Metal Rich Globular Clusters) The Faint Object Camera was used to make one observation of globular cluster NGC-6388 using special ultraviolet filters to look for metal rich stars. This will also help determine the stars temperature and luminosity. The observation was reported to have executed in fine lock without problems. Three observations were received and displayed. All three were reported to look normal. 1.22 Completed PC 4902 (HST Medium - Deep Survey - Part 2 low latitude Cycle 3 High) The Planetary Camera was used to make a long exposure taken in parallel with the prime science observation. Areas of interest are galactic structure and evolution. There is also the possibility of serendipitous discoveries. The observation was received and displayed in the OSS. It was reported to look okay. 1.23 Completed WFC 4383 (UV Spectroscopy of Two Very Carbon Poor Halo Planetary Nebulae - Cycle 3 Medium) The Wide Field Camera was used to make observations of the planetary nebula PK061+41.1 which is located in the halo of the Milky Way Galaxy. The observations was executed with no problems. The data was received and displayed. The image was reported to look normal. 1.24 Completed PC 4904 (The Nuclei of Nearby S0 and Spiral Galaxies - III A PC Snapshot Image) The Planetary Camera was used to make one observations of the low-red-shift galaxy NGC-2500 to study its surface brightness profile, identify sub-arc-second scale condensations, find point sources, and look for gasious disk and dust lanes. The observation was executed without problems. The data was received and displayed. The image was reported to look normal. 1.25 Completed HSP/UV2 3250 (Gravitational Lenes Part One - Cycle 2 High) The High Speed Photometer was used to make ultraviolet photometric observations of quasar QSO0957+56 which is an expected gravitational lenes system. It is hoped to find a characteristic signature in variability in the time scale of the order of weeks. Seven observations were received and displayed. They all were reported to look okay. 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 22 Successful : 22 Scheduled Reacquisitions : 23 Successful : 23 2.2 FHST Updates: Scheduled : 41 Successful : 41 2.3 Significant HSTARs: None 2.4 Operations Note: The JIS #4 (EVA #3) Simulation was successfully completed last Friday (Oct. 8, 1993) It was initiated at 1200Z, and lasted for 12 hours. All simulation objectives were met, and no major anomalies were reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC/PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 3.2 Complete FGS 5060 (Linking HIPPARCOS to the Extragalactic Reference Frame - Part 2 of 6) 3.3 Complete WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.5 Complete Two Sets of WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) 3.6 Complete GHRS/2 4067 (GHRS 1ST Order Grating Spectral Lamp Test) 3.7 Complete FGS 4893 (Orbital Motions of Close Binary Stars - cycle 3 High) 3.8 Complete WFC 4849 (GTO High Latitude Parallel Program - Cycle 3 High) 3.9 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.10 Complete FOS/RD 3791 (Quasar Absorption Line Survey: Cycle 2 Observations) 3.11 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 3.12 Complete Two sets of PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) 3.13 Complete Two Sets of HSP/PMT 3373 (The Size and Composition of Planetary Ring Particles - Cycle 2 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-